# Hardware Acceleration for Neural Networks: A Comprehensive Survey

### v1

@@ 2.3. Area and cost
Area and cost constraints shape what can be built and deployed at scale. Silicon area (in $\mathrm{mm}^{2}$) is consumed by compute arrays, on-chip SRAM, network-on-chip resources, and I/O PHYs, and each choice impacts achievable frequency and yield. Increasing on-chip memory can reduce off-chip traffic and improve performance-per-watt, but it increases die area and may limit clock rate; conversely, relying on off-chip memory shifts the bottleneck toward bandwidth and increases energy (Williams et al., 2009; Horowitz, 2014).
Figure 5. Trade-off between performance and area
@@ 2.3.4. FPGA-specific area/cost challenges
FPGAs trade silicon efficiency for reconfigurability; LUT/DSP/BRAM resources and routing overhead can limit peak throughput-per-area compared to ASICs (Nurvitadhi et al., 2017). Cost-efficiency depends on whether the deployment can keep the FPGA well utilized; otherwise fixed platform cost dominates. However, for deployments that value flexibility or low-volume customization, FPGAs can reduce overall engineering cost and time-to-deployment, particularly in cloud settings where the same fleet can be repurposed (Fowers et al., 2018). Toolchain productivity is also a cost factor: design iteration time and the ability to map new operators can dominate total cost of ownership (Venieris and Bouganis, 2017). Finally, partial reconfiguration and multi-tenant scheduling introduce additional complexity that can affect effective cost-per-inference (Fowers et al., 2018).
@@ 2.3.7. Neuromorphic area/cost challenges
Neuromorphic chips allocate area to distributed memory and communication fabrics to support event-driven execution. They can be cost-effective for always-on low-power sensing, but their value depends on workload fit and toolchain maturity rather than on conventional throughput-per-area metrics (Davies et al., 2018; Merolla et al., 2014). If the application requires significant preprocessing or dense computation, the effective cost rises because the neuromorphic chip must be paired with conventional processors (Gale et al., 2019). In addition, software ecosystem and deployment tooling influence engineering cost, which is often the limiting factor for adopting non-standard execution models (Hennessy and Patterson, 2019). As a result, neuromorphic cost-efficiency is best evaluated in complete systems and application pipelines, not in isolation (Coleman et al., 2019).
@@ 2.4. Performance limits from memory and communication
Figure 6. Memory and communication bottleneck
Figure 6 summarizes common memory and communication bottlenecks that dominate large-model training and serving at scale.
Distributed execution introduces collective communication and synchronization that can dominate iteration time. Data parallel training stresses all-reduce bandwidth, while tensor and pipeline parallelism introduce more frequent communication of activations and gradients; the optimal strategy depends on model size, sequence length, and interconnect characteristics (Shoeybi et al., 2019). Large-scale tensor-processor systems explicitly provision interconnect bandwidth and topology for these patterns, illustrating how system design shapes achievable scaling efficiency (Jouppi et al., 2021).
@@ 2.4.4. FPGA-specific memory/communication challenges
FPGA designs can minimize external traffic with streaming pipelines, but bandwidth to external DRAM and host-device I/O can dominate performance for large models or attention-heavy workloads (Nurvitadhi et al., 2017; Zhang et al., 2024a; Vaswani et al., 2017). For Transformers, buffering QKV projections and intermediate activations is challenging under limited on-chip BRAM, which can force frequent off-chip transfers and reduce throughput (Zhang et al., 2024a; Williams et al., 2009). In cloud deployments, PCIe/host interfaces and multi-tenant scheduling can become communication bottlenecks, and the end-to-end system often determines observed performance (Fowers et al., 2018). Toolchain decisions (HLS vs. RTL, memory partitioning) affect achievable bandwidth and hence the effective communication bottleneck (Venieris and Bouganis, 2017).
@@ 2.4.7. Neuromorphic memory/communication challenges
Neuromorphic systems distribute memory and communication across cores; performance depends on sparse event traffic and locality (Davies et al., 2018; Merolla et al., 2014). When spiking activity is dense, communication fabric contention can dominate, and the energy/latency benefits of event-driven execution diminish (Davies et al., 2018). End-to-end communication overhead also includes encoding/decoding between conventional sensors and spiking representations, which can outweigh benefits if sparsity is not intrinsic (Gale et al., 2019). These effects mean that neuromorphic communication challenges are best evaluated at the application level rather than by peak metrics alone (Coleman et al., 2019).
@@ 2.5. Resource utilization and contention
Resource consumption is broader than power and area: it includes compute utilization, memory footprint, bandwidth demand, and contention for shared resources. Real systems are constrained by multiple shared bottlenecks (SM occupancy, cache capacity, memory controllers, and interconnect), so maximizing one resource in isolation can degrade overall throughput by increasing contention elsewhere (Williams et al., 2009).
@@
Figure 7. Resource utilization

### v3

@@ 2.2.1. GPU-specific throughput/latency challenges
Achieving consistent p99 latency on GPUs is further complicated by the “bulk synchronous” programming model: a single long-running kernel can block other streams, making it hard to multiplex small, latency-critical tasks with background throughput-oriented work. Context switching on GPUs is relatively expensive, so preemption strategies for real-time serving are limited compared to CPUs. This forces serving systems to rely on cooperative scheduling (e.g., careful batching windows) or multi-instance GPU (MIG) partitioning to isolate workloads, though MIG partitions compute and memory statically rather than dynamically (NVIDIA, 2023; Williams et al., 2009).
@@ 2.2.2. TPU/NPU-specific throughput/latency challenges
A key trade-off for TPU/NPU architectures is the “compile-time vs. run-time” balance. Relying on static compilation provides excellent determinism and dense packing for fixed graphs, but it can lead to severe latency penalties if dynamic shapes force recompilation or padding. For variable-length sequences (common in LLM serving), the system must either pad to the worst case (wasting throughput) or use specialized dynamic-shape compilers that may generate less optimal code. Managing this trade-off requires sophisticated runtime systems that can bucket requests by size or switch between pre-compiled binaries on the fly, adding complexity to the serving stack (Jouppi et al., 2021; Kwon et al., 2023).
@@ 2.2.3. ASIC-specific throughput/latency challenges
Furthermore, the “throughput at what cost” question is acute for ASICs: maximizing peak Tera Operations Per Second (TOPS) often leads to large systolic arrays that have high startup latency and poor efficiency for small batches (batch-1 inference). To target low-latency serving, designers may split resources into multiple smaller cores or use finer-grained pipelining, but this complicates synchronization and on-chip interconnect design. If the workload is memory-bound (e.g., decode phase of LLMs), the ASIC’s peak compute advantage becomes irrelevant, and latency is governed solely by the available HBM bandwidth and the efficiency of the memory controller (Groq, 2024; Williams et al., 2009).
@@ 2.2.4. FPGA-specific throughput/latency challenges
However, achieving this determinism requires resolving the “routing vs. logic” frequency wall. Deeply pipelined designs can run at high clock speeds, but complex control logic or irregular memory access patterns (e.g., from sparse attention) can create routing congestion that lowers the achievable frequency, directly hurting throughput. Moreover, if the model does not fit entirely on-chip, the FPGA must buffer and swap weights/activations; if this swapping is not perfectly hidden by compute (double buffering), latency spikes occur. Modern FPGAs with HBM stacks help alleviate the bandwidth bottleneck, but maximizing their effective bandwidth requires wide data paths and careful memory controller tuning, which complicates the High-Level Synthesis (HLS)/Register Transfer Level (RTL) design effort (Nurvitadhi et al., 2017; Li et al., 2024a).
@@ 2.2.5. LPU/LLM-serving throughput/latency challenges
Ideally, serving systems want throughput without latency degradation, but batching inherently trades one for the other. Large batches improve GEMM efficiency and aggregate bandwidth utilization, but they increase the time each token spends waiting in queues or for other sequences to finish their step. This “batching capability gap” is where specialized architectures try to innovate: by managing control flow and dependency tracking in hardware (rather than GPU kernels + CPU driver), LPUs attempt to make fine-grained interleaved execution feasible, allowing high concurrency with lower queuing delays. Even so, the physical reality of moving KV-cache bytes means that scaling context length inevitably hurts latency unless memory bandwidth scales proportionally (Groq, 2024; Williams et al., 2009).
@@ 2.2.6. In-/near-memory and analog throughput/latency challenges
A deeper latency challenge stems from the dataflow between analog and digital domains. If a network requires frequent normalization or activation functions that are best done digitally, data must be repeatedly converted and moved, killing the latency benefits of in-place compute. Pipelining these stages can hide throughput costs but increases pipeline depth (latency). Furthermore, writing weights to analog arrays (e.g., Resistive RAM (ReRAM) or Flash) is often slow and energy-intensive; this makes analog approaches less suitable for workloads requiring rapid model switching or dynamic LoRA adapters, restricting their “speed” advantage to static-weight inference scenarios (Shafiee et al., 2016; Kim et al., 2025).
@@ 2.2.7. Neuromorphic throughput/latency challenges
Throughput in neuromorphic systems is also non-standard: it is often measured in “events per second” or “inferences per second” under a specific sparsity assumption. If the input data becomes dense (e.g., a complex visual scene), the event traffic can saturate the on-chip interconnect, causing spikes to be dropped or delayed, which degrades both accuracy and latency. This traffic-dependent latency profile makes it hard to guarantee real-time bounds compared to the predictable scan-line processing of a traditional accelerator. Thus, achieving high speed requires not just fast silicon but also algorithms that actively maintain sparsity and locality throughout the network execution (Davies et al., 2018; Coleman et al., 2017).
@@ 2.3. Area and cost
Area and cost constraints shape what can be built and deployed at scale. Silicon area (in $\mathrm{mm}^{2}$) is consumed by compute arrays, on-chip SRAM, Network-on-Chip (NoC) resources, and I/O Physical Layers (PHYs), and each choice impacts achievable frequency and yield. Increasing on-chip memory can reduce off-chip traffic and improve performance-per-watt, but it increases die area and may limit clock rate; conversely, relying on off-chip memory shifts the bottleneck toward bandwidth and increases energy (Williams et al., 2009; Horowitz, 2014).
Figure 5. The fundamental trade-off between performance and silicon area/cost, illustrating how increasing parallelism and memory capacity improves throughput but raises manufacturing and packaging costs.
@@ 2.3.1. GPU-specific area/cost challenges
A major driver of GPU cost is the memory subsystem: HBM stacks and CoWoS (Chip-on-Wafer-on-Substrate) packaging are significant fractions of the BOM (Bill of Materials). For inference workloads that are memory-capacity bound (like hosting LLaMA-70B), users must purchase more GPUs just to fit the model weights and KV-cache, even if compute utilization is low. This “capacity tax” has spurred interest in memory-expansion technologies (e.g., CXL) and aggressive quantization (4-bit/3-bit), which effectively reduce the dollar-cost per parameter by fitting larger models onto fewer devices. However, these solutions trade off bandwidth or accuracy, reinforcing the link between algorithmic choices and deployment economics (Micikevicius et al., 2018; Kwon et al., 2023).
@@ 2.3.2. TPU/NPU-specific area/cost challenges
For datacenter TPUs, the “pod” architecture amortizes the cost of the high-speed torus interconnect across many chips, but this makes the system expensive to deploy in small increments. The fixed topology is highly efficient for large-scale training jobs that map well to the torus, but it can be rigid for fragmented inference workloads or for models that don’t naturally shard across the pod’s dimensions. This creates a utilization challenge: if a job doesn’t need a full pod slice, resources may be stranded. Thus, the total cost of ownership (TCO) depends on the scheduler’s ability to bin-pack diverse jobs onto the available topology without fragmentation (Jouppi et al., 2021).
@@ 2.3.3. ASIC-specific area/cost challenges
Furthermore, the “dark silicon” problem manifests in ASICs when new model features emerge that the hardware wasn’t designed for. If an accelerator devotes 30% of its area to a specific sparsity engine or activation function that later falls out of favor (e.g., a move from ReLU to Swish/GELU), that area becomes wasted cost. This risk forces architects to over-provision programmable elements (vector processors or DSPs) alongside fixed-function units, diluting the area-efficiency advantage. As a result, the most successful commercial ASICs often pair efficient matrix cores with a fairly capable general-purpose vector Instruction Set Architecture (ISA) to future-proof the investment (Hennessy and Patterson, 2019; Williams et al., 2009).
@@ 2.3.4. FPGA-specific area/cost challenges
FPGAs trade silicon efficiency for reconfigurability; Look-Up Table (LUT)/DSP/BRAM resources and routing overhead can limit peak throughput-per-area compared to ASICs (Nurvitadhi et al., 2017). Cost-efficiency depends on whether the deployment can keep the FPGA well utilized; otherwise fixed platform cost dominates. However, for deployments that value flexibility or low-volume customization, FPGAs can reduce overall engineering cost and time-to-deployment, particularly in cloud settings where the same fleet can be repurposed (Fowers et al., 2018). Toolchain productivity is also a cost factor: design iteration time and the ability to map new operators can dominate total cost of ownership (Venieris and Bouganis, 2016). Finally, partial reconfiguration and multi-tenant scheduling introduce additional complexity that can affect effective cost-per-inference (Fowers et al., 2018).
High-end FPGAs with HBM and hardened NoCs are expensive, often rivaling GPUs in unit cost. To justify this premium, the design must deliver value that GPUs cannot—typically strictly deterministic latency or direct integration with network/storage (“smart Network Interface Card (NIC)” or “smart Solid State Drive (SSD)” use cases). If the FPGA is used merely as a slower GPU for batch processing, the area/cost math rarely works out due to the overhead of programmable fabric. Therefore, cost-effective FPGA deployment often involves moving the compute to the data (e.g., inside the network switch) to save system-level energy and bandwidth, rather than competing on raw FLOPs/\$ in a rack server (Fowers et al., 2018; Nurvitadhi et al., 2017).
@@ 2.3.5. LPU/LLM-serving area/cost challenges
A specific cost driver for LPU serving is the need for SRAM-heavy architectures to hide latency. SRAM is orders of magnitude more expensive per bit than DRAM. While an SRAM-centric design (like Groq’s) provides unmatched deterministic latency and throughput for small batches, scaling it to store the weights of a 70B+ parameter model requires daisy-chaining many chips, exploding the system cost. This creates a bifurcation in the market: SRAM-based LPUs for latency-critical, lower-parameter (or sharded) serving, and HBM-based GPUs/ASICs for cost-optimized, high-capacity serving. The “right” cost choice depends entirely on the user’s willingness to pay for milliseconds of latency reduction (Groq, 2024; Williams et al., 2009).
@@ 2.3.6. In-/near-memory and analog area/cost challenges
Fabrication cost is another hurdle: many dense non-volatile memory technologies (Resistive RAM (RRAM), Phase-Change Memory (PCM)) require specialized process steps that may not be available in standard logic finFET nodes, or they require integration via expensive 2.5D/3D stacking. This splits the manufacturing ecosystem and can delay access to the latest lithography nodes that benefit the digital control logic. Consequently, analog accelerators often lag behind digital counterparts in frequency and logic density, forcing them to compete purely on energy efficiency in niche markets (e.g., edge, IoT) rather than replacing general-purpose datacenter silicon (Hennessy and Patterson, 2019; Chi et al., 2016).
@@ 2.3.7. Neuromorphic area/cost challenges
Neuromorphic chips allocate area to distributed memory and communication fabrics to support event-driven execution. They can be cost-effective for always-on low-power sensing, but their value depends on workload fit and toolchain maturity rather than on conventional throughput-per-area metrics (Davies et al., 2018; Merolla et al., 2014). If the application requires significant preprocessing or dense computation, the effective cost rises because the neuromorphic chip must be paired with conventional processors (Gale et al., 2019). In addition, software ecosystem and deployment tooling influence engineering cost, which is often the limiting factor for adopting non-standard execution models (Hennessy and Patterson, 2019). As a result, neuromorphic cost-efficiency is best evaluated in complete systems and application pipelines, not in isolation (Coleman et al., 2017).
The area trade-off in neuromorphic designs is unique: they sacrifice dense arithmetic logic (large MAC arrays) to prioritize state storage (synapses) and fine-grained routing. For sparse workloads, this is area-efficient because inactive circuits consume little static power and no dynamic power. However, for dense workloads, the lack of time-multiplexed arithmetic units means the chip must physically instantiate more neurons/synapses to represent the model, potentially leading to a larger die size than a compact, reused systolic array. Thus, the “cost” of neuromorphic hardware is only justified when the sparsity factor is high enough to offset the lower area-density of computation (Davies et al., 2018).
@@ 2.4. Performance limits from memory and communication
Figure 6. Memory and communication bottlenecks in large-scale training and inference, emphasizing the impact of limited HBM bandwidth, interconnect latency, and the overhead of collective communication primitives.
Figure 6 summarizes common memory and communication bottlenecks that dominate large-model training and serving at scale. Distributed execution introduces collective communication and synchronization that can dominate iteration time. Data parallel training stresses all-reduce bandwidth, while tensor and pipeline parallelism introduce more frequent communication of activations and gradients; the optimal strategy depends on model size, sequence length, and interconnect characteristics (Shoeybi et al., 2019). Large-scale tensor-processor systems explicitly provision interconnect bandwidth and topology for these patterns, illustrating how system design shapes achievable scaling efficiency (Jouppi et al., 2021).
@@ 2.4.1. GPU-specific memory/communication challenges
Specifically, the “memory wall” on GPUs is exacerbated by the growing disparity between arithmetic throughput (scaling rapidly with tensor cores) and memory bandwidth (scaling more slowly). This puts pressure on software to maximize cache reuse through techniques like activation checkpointing (trading compute for memory) and aggressive kernel fusion (keeping data in registers). In distributed settings, communication bandwidth (NVLink/Infiniband) often determines the feasibility of splitting a single model across devices. If the interconnect is slow, tensor parallelism becomes prohibitively expensive due to frequent all-reduce operations, forcing users into less efficient pipeline parallelism or pure data parallelism with lower effective batch sizes (Shoeybi et al., 2019).
@@ 2.4.2. TPU/NPU-specific memory/communication challenges
A distinct challenge for TPUs is the management of “HBM fragmentation” and scratchpad allocation. The compiler must orchestrate data movement between HBM and on-chip memory with cycle-level precision. If a model’s activation working set slightly exceeds the SRAM capacity, the compiler may spill data to HBM, causing a sharp “performance cliff.” This binary behavior (fit vs. spill) makes performance less predictable than on cache-based architectures (GPUs), where performance degrades more gracefully. Consequently, developers often spend significant effort manual tuning batch sizes and partition strategies to stay on the “fast path” of the memory hierarchy (Jouppi et al., 2017; Williams et al., 2009).
@@ 2.4.3. ASIC-specific memory/communication challenges
The “bandwidth tax” of programmability also affects ASICs. If an architecture relies on a host CPU to issue commands for every tile or layer, the PCIe bus or command-queue latency can become a bottleneck, especially for small-batch inference. To avoid this, advanced ASICs implement autonomous command processors or graph executors on-chip, allowing the device to run entire subgraphs without host intervention. However, this increases design complexity and requires a robust compiler to generate the command streams. Without this autonomy, the accelerator may sit idle waiting for the host to pointer-chase through the next set of descriptors, wasting the available DRAM bandwidth (Hennessy and Patterson, 2019).
@@ 2.4.4. FPGA-specific memory/communication challenges
FPGA designs can minimize external traffic with streaming pipelines, but bandwidth to external DRAM and host-device I/O can dominate performance for large models or attention-heavy workloads (Nurvitadhi et al., 2017; Zhang et al., 2024a; Vaswani et al., 2017). For Transformers, buffering Query-Key-Value (QKV) projections and intermediate activations is challenging under limited on-chip BRAM, which can force frequent off-chip transfers and reduce throughput (Zhang et al., 2024a; Williams et al., 2009). In cloud deployments, PCIe/host interfaces and multi-tenant scheduling can become communication bottlenecks, and the end-to-end system often determines observed performance (Fowers et al., 2018). Toolchain decisions (HLS vs. RTL, memory partitioning) affect achievable bandwidth and hence the effective communication bottleneck (Venieris and Bouganis, 2016).
Another constraint is the “port limitation” of on-chip memory. BRAMs typically have only two ports. If a compute kernel needs to read 16 operands per cycle to keep the pipeline full, the designer must bank and duplicate memory extensively or run the BRAMs at a higher clock multiple (if possible). This memory partitioning problem is NP-hard and is a frequent cause of routing congestion or lower-than-peak performance. HLS tools try to automate this, but for complex access patterns (like sliding windows with non-unit strides or sparse lookups), manual pragma insertion or RTL rewriting is often needed to saturate the external memory interface (Zhang et al., 2015; Canis et al., 2013).
@@ 2.4.5. LPU/LLM-serving memory/communication challenges
Communication between chips is critical for LPUs because a single chip rarely holds a full large model. The interconnect must support extremely low-latency, fine-grained tensor slicing to allow “tensor-parallel” execution across 16–64 chips without stall bubbles. Unlike GPU clusters that might tolerate microseconds of all-reduce latency, an LPU pipeline executing layer-by-layer requires nanosecond-scale synchronization to maintain its deterministic throughput guarantees. This drives LPU interconnects to be proprietary, statically routed, and integrated directly into the ISA, rather than relying on standard Ethernet or PCIe switches (Groq, 2024; Shazeer et al., 2017).
@@ 2.4.6. In-/near-memory and analog memory/communication challenges
Additionally, the “fan-out/fan-in” communication within a crossbar array poses signal integrity challenges. As arrays grow larger to hold more weights, the analog signal path lengthens (increasing IR drop and capacitance), which limits the readout speed or precision. Digital communication between tiles (e.g., aggregating partial sums from multiple arrays) then becomes the new bottleneck. If the NoC connecting these tiles isn’t provisioned with enough bandwidth, the fast analog cores will stall waiting to retire their results, negating the throughput benefit. Thus, PIM architectures effectively trade off-chip memory bandwidth problems for on-chip network design challenges (Shafiee et al., 2016; Kim et al., 2025).
@@ 2.4.7. Neuromorphic memory/communication challenges
Neuromorphic systems distribute memory and communication across cores; performance depends on sparse event traffic and locality (Davies et al., 2018; Merolla et al., 2014). When spiking activity is dense, communication fabric contention can dominate, and the energy/latency benefits of event-driven execution diminish (Davies et al., 2018). End-to-end communication overhead also includes encoding/decoding between conventional sensors and spiking representations, which can outweigh benefits if sparsity is not intrinsic (Gale et al., 2019). These effects mean that neuromorphic communication challenges are best evaluated at the application level rather than by peak metrics alone (Coleman et al., 2017).
Scaling neuromorphic systems to large models introduces a “synaptic storage” challenge. While distributed SRAM is fast, it is not as dense as DRAM. Storing a billion-parameter model entirely in on-chip SRAM is cost-prohibitive. Some designs use multi-chip meshes, but this reintroduces inter-chip communication latency, which can disrupt the precise timing required for spike-based plasticity (Spike-Timing-Dependent Plasticity (STDP)). Therefore, efficient memory virtualization or hierarchical storage (local SRAM + backing DRAM) with smart prefetching is needed to support scale-out without killing the event-driven efficiency (Davies et al., 2018).
@@ 2.5. Resource utilization and contention
Resource consumption is broader than power and area: it includes compute utilization, memory footprint, bandwidth demand, and contention for shared resources. Real systems are constrained by multiple shared bottlenecks (Streaming Multiprocessor (SM) occupancy, cache capacity, memory controllers, and interconnect), so maximizing one resource in isolation can degrade overall throughput by increasing contention elsewhere (Williams et al., 2009).
@@
Figure 7. Resource utilization challenges arising from irregular workloads, showing how sparsity, dynamic shapes, and conditional execution (e.g., MoE) can lead to load imbalance and underutilized compute units.
@@ 2.5.1. GPU-specific utilization challenges
Another utilization pitfall is the “tail effect” in distributed training. If one GPU in a 1000-GPU cluster is slow (due to thermal throttling, Error Correction Code (ECC) errors, or straggler tasks), the entire synchronous training step waits, effectively reducing the cluster-wide utilization to zero during the wait time. Fault tolerance and straggler mitigation strategies are thus essential for utilization at scale. Similarly, in pipeline parallelism, “pipeline bubbles” (idle time during fill/drain phases) reduce effective FLOPs. Minimizing these bubbles requires complex interleaved schedules (e.g., 1F1B) that increase memory pressure, creating a tension between utilization and memory capacity (Shoeybi et al., 2019; Rajbhandari et al., 2020).